The distance around a tennis court is 330 feet. If the length of the court is 9 feet more than twice the width, find the dimensions of the tennis court.
Answer by stanbon(75887)

The distance around a tennis court is 330 feet. If the length of the court is 9 feet more than twice the width, find the dimensions of the tennis court.
---------
Perimeter = 2(length + width)
330 = 2(2x+9 + x)
165 = 3x+9
3x = 156
x = 52 (width)
-----
length = 2x+9 = 2*52+9 = 113
==================================
Cheers,
Stan H.
==============